OpenStudy (anonymous):

How would I find the instantaneous rate of change using this formula y=3.9657(0.9982^x) and given a table of values?

OpenStudy (anonymous):

789

OpenStudy (anonymous):

greta knows is cursing at me

OpenStudy (p0sitr0n):

you could proceed multiple ways, but one would be to differentiate y with respect to x, and then plug in some value for x

OpenStudy (anonymous):

ok thanks! how would I go about differentiating the function?

OpenStudy (alekos):

\[d/dx(a b^{x}) = (alnb)b^{x} \]

OpenStudy (anonymous):

Oh so: y=3.9657(0.9982^x) y'=3.9657(0.9982^x) ln0.9982 (if x=30) y'=3.9657(0.9982^(30)) ln0.9982 y'=-0.00677